OpenStudy (ibbutibbu.):

Can someone help me? |9x + 4| - 11 = -4

OpenStudy (anonymous):

First, isolate the absolute value function by adding 11 to both sides. What do you get?

OpenStudy (hba):

Follow those steps: 1) add 11 on both sides 2) subtract 4 on both sides 3) Divide both sides by 9

OpenStudy (ibbutibbu.):

|9x +15| = -15 ?

OpenStudy (freckles):

@hba that will only give one solution this should give two since we will have |f|=a where a is positive |f|=a means f=a or f=-a

OpenStudy (anonymous):

Add 11, not subtract

OpenStudy (freckles):

well 11-4 isn't -15

OpenStudy (ibbutibbu.):

oh lol i did -11-4 so its actually 7

OpenStudy (anonymous):

Great now you have\[\left| 9x+4 \right| = 7\]One solution is when the quantity (9x+4) is positive, so solve\[9x+4=7\]The other solution is when the quantity (9x+4) is negative. So, solve\[-\left( 9x+4 \right)=7\]Then you'll have the complete solution

OpenStudy (ibbutibbu.):

so the first one is x = 1/3 and the 2nd one is x = -11/9 ?

OpenStudy (anonymous):

Excellent. Well done!

OpenStudy (ibbutibbu.):

yay thx

OpenStudy (anonymous):

you're welcome